OpenStudy (anonymous):

{x+4y-z=3 {x+5y-2z=5 {3x+12y-3z=9 possible answers: x = 3z + 5, and y = z - 2, where z is any real number or {(x, y, z) |x = 3z + 5, and y = z - 2, where z is any real number} x = z - 2, and y = -3z - 5, where z is any real number or {(x, y, z) |x = z - 2, and y = -3z - 5, where z is any real number} x = -3z - 5, and y = z + 2, where z is any real number or {(x, y, z) |x = -3z - 5, and y = z + 2, where z is any real number} inconsistent

OpenStudy (mertsj):

-2x-8y+2z=-6 x+5y-2x=5 Adding: -x-3y=-1

OpenStudy (mertsj):

-3x-12y+3z=-9 3x+12y-3z=9 0=0

OpenStudy (anonymous):

so whats the answer ? you are telling me that this is dependent?

OpenStudy (mertsj):

x=-3z-5 y=z+2

OpenStudy (mertsj):

The third choice.

OpenStudy (anonymous):

thank you sir I really appreciate it

OpenStudy (mertsj):

yw